# Distributional Evaluation of Generative Models via Relative Density Ratio

### v1

@@ 4.2 CelebA-64
Figure 5: CelebA-64: DDIM evaluated on full training data.
DDIM. Denoising diffusion models such as DDIM (Song et al., 2021b) are known to generate highly realistic images, but the sampling process is relatively slow. Therefore, in this experiment, when evaluating the pretrained DDIM model (Song et al., 2021a), we first generate 160,000 images and store them on disk. We then train the ratio function stochastically using batch updates from this pre-generated dataset. Figure 5 provides the analysis result on the DDIM generator. Even though the DDIM generated images in Figure 5(e) are visually realistic, but the vast distributional discrepancy shown in Figure 5(a) and Table 5(b) suggests otherwise. The DDIM generator has good coverage, since $r(X_{p})$ and $r(X_{q})$ have a similar range from near 0 to near 2. However, a vast majority of the DDIM-generated samples are highly concentrated around $r(X_{q})$ near 0. In fact, 157,779 out of 160,000 generated DDIM samples have $r(X_{q})$ values smaller than the least $r(X_{p})=0.00012$, completely out of the support of the real data. The left panel in Figure 5(e) shows such examples (more such examples in Supplementary Figure 7), and they are all centered, frontal or near-frontal portraits, with neutral to positive expressions, simple backgrounds, and the lighting appears relatively even and bright, without strong shadows. The overall structure of these faces is consistent, suggesting DDIM generated them from a tight distribution around a latent region representing a certain geometric average of the training data. To further study and interpret such distributional differences, we run a logistic regression on the real training data on $P(r(X_{p})>1)$, and use the 40 attributes as the covariates. Table 5(b) further reveals some key features ranked by the smallest p-values of this difference: for example, the significant attribute with the largest effect is the 5_o_Clock_Shadow, meaning that it is difficult for DDIM to generate such images with shadows; on the other hand, it is easy for DDIM to generate realistic images with people wearing hat. These identified features can also be visually checked by comparing the middle and right panels in Figure 5(d) and Figure 5(e).
4.3 American Gut Project
For this analysis, we use the 2018 release (McDonald et al., 2018a) of the American Gut Project (McDonald et al., 2018c) data, and the preprocessing is publicly available in our GitHub page (https://github.com/yuliangxu/RDR). We use the relative abundance data derived from the taxa count data, such that the sum over all taxa for one sample equals 1. After preprocessing, we have 614 taxa (columns) at the species level, and split the data into 80% for training and 20% for testing, resulting in training data of $n=10443$ and test data $n=2611$. We apply the ICFM (Independent Continuous Flow Matching) (Lipman et al., 2022; Tong et al., 2023) as the generator on the logistic tree transformed space, where the relative abundance is aggregated through all levels in the phylogenetic tree and mapped to the real line using logit transform.
(a) Histogram.
$r(X_{p})$ $r(X_{q})$
length 2611 2611
mean 1.91 0.02
std 0.13 0.20
min 0.78 0.00
q1 1.91 0.00
median 1.96 0.00
q3 1.98 0.00
max 2.00 1.99
(b) Summary Statistics.
(c) PCoA plot with Bray-Curtis metric.
(d) Stacked bar plot comparison on relative abundance.
Figure 6: American Gut: ICFM evaluated on testing data.
Figure 6(d) shows the stacked bar plot in the simplex space of the generated sample (top) and the real sample (bottom). Each column in the stacked bar plot represents one sample, and the y-axis is the proportional breakdown in each class-level relative abundance for better visualization. The samples are ordered by the largest abundance of taxa. Different from the previous CelebA example, where we can visually check the difference between the generated sample and the real sample, visually evaluating the generation quality of the relative abundance data becomes impossible, even with a much smaller dimensionality. Some common visualizations include the alpha diversities to compare within-sample diversity (Figure 8) and beta diversities such as the PCoA scatter plot using Bray-Curtis metric (Figure 6(c)) to compare between-sample diversity. The ICFM generator performs well in all of these visual comparisons.
To have a deeper understanding of this distributional difference, in particular, which taxa are over-represented/under-represented in the generated sample, we conduct an association test between each level of taxa abundance and the learned ratio. The generator and the ratio learn are both trained on the lowest species level. For the association test, we first transform the relative abundance samples using the central log ratio (CLR) transformation, then stack the CLR-transformed real and generated samples, as well as the learned ratio for both samples. At each taxonomic level, we aggregate the relative abundance to that level, and run a Spearman correlation test between the learned ratio and the CLR-transformed samples. Figure 9 shows the correlation coefficients at each taxonomic level. If the correlation coefficient for taxon $k$ is positive (red), this means that when $r(X)$ is higher, taxon $k$ increases relative to the geometric mean of its level, and is more present in the real sample than the generated sample. In Figure 9, the size of each node (circle sector) reflects the number of descendant taxa contained within that lineage. Based on Figure 9, we can see the tendency that the ICFM-generated sample tends to over-represent taxa with few descendants, whereas for large classes such as Clostridia, ICFM tends to have less representation (this is also hinted by Figure 6(c)). In a way, similar to the CelebA result, ICFM as a generator tends to concentrate its mass towards a more averaged area in the simplex.
5 Discussion
Acknowledgement
We are grateful to Naoki Awaya for his valuable comments and discussions. We thank Ganchao Wei for providing the code to preprocess the American Gut Data and the code to train the ICFM generator on the relative abundance data. This work is supported by NIGMS grant R01-GM135440.
Data Availability
Our GitHub page (https://github.com/yuliangxu/RDR) contains all the reproducible workflows to recreate all the analysis results reported in this manuscript. For the 1D examples and MNIST examples, we provide the Jupyter notebook with the same plots used in the manuscript. For the CelebA-64 example, we provide the pipeline to generate 160,000 training images, and the workflow to recreate the analysis results given the pre-generated data. For the American Gut example, we provide the data preprocessing code and the workflow for the main result in the paper. All data used are openly available, and we have included the code to directly download the data online. We use pre-trained models for MNIST and CelebA-64 examples, and have included the URL to download the pretrained checkpoints in our reproducible workflow.
Appendix A Proof of Proposition 1
Proof.
Appendix B Proofs for general estimators
B.1 Proof of Theorem 1
Proof.
According to the relative data processing inequality, $D_{\phi}(P\|Q)\geq D_{\phi}(F_{\#}P\|F_{\#}Q)$. By Theorem 3.3 in Liese (2012) equivalent conditions (A) and (C), the equality holds when the transformation $F$ is any sufficient statistic for the family of probabilities $\mathcal{P}:=\{P,Q\}$.
According to Theorem 6.12 in Chapter 1 in Lehmann and Casella (2006) (which requires $P$ and $Q$ have common support), the density ratio $g(X)=\frac{p(X)}{q(X)}$ is a minimal sufficient statistic for the family $\mathcal{P}$. Hence $g$ is a maximizer for $D_{\phi}(F_{\#}P\|F_{\#}Q)$ among all other measurable maps $F:\mathcal{X}\to\mathcal{Y}$.
B.2 Proof of Theorem 2
Proof.
$\left|\int g^{-1/2}\text{d}(P_{n}-P)+\int g^{1/2}\text{d}(Q_{m}-Q)\right|$
By Assumption 3 and Assumption 4, and Theorem 3.7 in van de Geer (2000),
Next, to show the consistency of $\hat{g}_{n}$, we first define
To show the lower bound, note that for any $g\in\mathcal{G}$,
$=\int\frac{\left(\sqrt{g}-\sqrt{g_{0}}\right)^{2}}{\sqrt{g}}\text{d}Q.$
B.3 Proof of Lemma 2
Proof.
This result is mentioned in Nguyen et al. (2010). We give a formal proof here. By Theorem 2.4.1 in van der Vaart and Wellner (1996), the bracketing number $N_{[]}(\epsilon,\mathcal{F},L^{1}(P))<\infty$ implies $\mathcal{F}$ is Glivenko-Cantelli, i.e. $\sup_{f\in\mathcal{F}}|P_{n}f-Pf|\to 0$ a.s.
B.4 Proof of Theorem 3
Proof.
This proof follows Theorem 3.2.5 in van der Vaart and Wellner (1996).
Now we need to verify the following four conditions from Theorem 3.2.5 in van der Vaart and Wellner (1996).
Conditions to be verified:
  1. 1.
    $\mathbb{M}(g)-\mathbb{M}(g_{0})\lesssim-h_{Q}^{2}(g,g_{0})$;
  2. 2.
    For every $n$ and sufficiently small $\delta$,
  3. 3.
  4. 4.
    $h_{Q}(\hat{g}_{n},g_{0})\overset{p}{\to}0$. [^1] The original Theorem 3.2.5 in van der Vaart and Wellner (1996) is stated with outer probability in Condition 4 and outer expectation in Condition 2. Here, we omit the discussion on the measurability of $\hat{g}_{n}$.
We adopt the following empirical process notations
Hence we only need to find $\phi_{n}(\delta)$ such that
We use Lemma 3.4.2 in van der Vaart and Wellner (1996) to show the entropy condition. We first show (B), and the proof for (A) is similar. By the definition (8), for any $f\in\mathcal{F}_{\delta}^{Q}$, $Qf^{2}<\delta^{2}$, and Assumption 5 implies there exists a positive constant $M_{\mathcal{F}^{Q}}$ such that $\sup_{g\in\mathcal{G}}|g^{1/2}-g_{0}^{1/2}|\leq M_{\mathcal{F}^{Q}}<\infty$. By Lemma 3.4.2 in van der Vaart and Wellner (1996)
According to Assumption 6, for any small $\delta<\eta$,
Appendix C Proofs for neural network sieved space
To show the rate of the estimation error, we need to verify conditions in Theorem 3.4.1 in van der Vaart and Wellner (1996).
Conditions to be verified:
For every $n$ and $\delta_{n}<\delta<\infty$,
  1. 1.
  2. 2.
  3. 3.
  4. 4.
With these conditions, one can conclude that $r_{n}h_{Q}(\hat{g}_{n},g_{n})=O_{p}(1)$. Here, $\delta_{n}$ is a multiple of $h_{Q}(g_{n},g_{0})$. Note that $h_{Q}(g_{n},g_{0})$ depends on the approximation rate in Proposition 2 of Suzuki and Nitanda (2021). The sketch of the proof is as follows: first, we use Lemma 3 to show Condition 1; then use Lemma 4 to compute the bracketing number of the sparse neural network space that will be used in the proof of Condition 2; finally, we will show Condition 2 to 4 in the proof of Theorem 4, and solve for the rate $r_{n}$ in Condition 3 where $r_{n}\asymp\delta_{n}^{-1}$, hence the final convergence rate will be the same order of $h_{Q}(\hat{g}_{n},g_{n})$ and $h_{Q}(g_{n},g_{0})$.
Lemma 3.
Proof.
The intuition behind this proof is that, for $g_{n}$ in $\mathcal{G}_{n}$, we want $g_{n}$ to be closer to the global optimizer $g_{0}$ than the arbitrary $g$ when verifying the first condition in Theorem 3.4.1 in van der Vaart and Wellner (1996).
For the objective $\mathbb{M}(g)$ defined in (17),
For the first term, based on Assumption 5,
$\leq 2M_{l}\sqrt{M_{u}}h_{Q}(g_{0},g_{n})h_{Q}(g,g_{n}).$
$\mathbb{M}(g)-\mathbb{M}(g_{n})$ $\leq-\frac{1}{2\sqrt{M_{u}}}h^{2}_{Q}(g,g_{n}).$
Lemma 4.
Proof.
This proof partially follows the ideas in Lemma S.2 in Kaji et al. (2023) and Lemma 6 in Suzuki and Nitanda (2021). The key step is to verify the conditions for Theorem 2.7.11 in van der Vaart and Wellner (1996).
$\leq UB\left\|f^{(l-1)}\circ\dots\circ f^{(1)}\right\|_{\infty}+B$
$=U^{l-1}B^{l}(d\|x\|_{\infty}+2).$
$\leq$ $(UB)^{L-1}H(x)d_{w}^{(L)}+d_{b}^{(L)}+$
[…]
Denote $F(x):=L(UB)^{L-1}H(x)=L(UB)^{L-1}\left(d\|x\|_{\infty}+2\right)$. Also, note that using the vector infinity norm and radius $\epsilon$, the covering number of $B,B]^{S}$ is $(B/\epsilon)^{S}$. According to Theorem 2.7.11 in van der Vaart and Wellner (1996), then
Hence,
C.1 Proof of Theorem 4
Proof.
By Lemma 3.4.2 in van der Vaart and Wellner (1996), if there exists $M$ such that $\|g_{n}^{1/2}-g^{1/2}\|_{\infty}\leq M$, and $Q\left\{\left(g^{1/2}-g_{n}^{1/2}\right)^{2}\right\}\lesssim\delta^{2}$, then
and note that
hence
For the upper bound of $\mathbb{E}\|\mathbb{G}_{n}^{P}\|_{\mathcal{F}_{\delta,n}^{P}}$, we can similarly apply Lemma 3.4.5 in van der Vaart and Wellner (1996). To verify that $P\left\{\left(g^{-1/2}-g_{n}^{-1/2}\right)^{2}\right\}\lesssim\delta^{2}$ given that in $\mathcal{F}_{\delta,n}^{P}$, $h_{Q}(g_{n},g)<\delta$, we refer to the previous proof in (21) and replace $g$ by $g_{n}$ in (21). Similarly, $\left\|\frac{1}{\sqrt{g^{L}}}-\frac{1}{\sqrt{g^{U}}}\right\|_{L^{2}(P)}=\left\|\frac{g^{L}-g^{U}}{\sqrt{g^{U}g^{L}}\left(\sqrt{g^{U}}+\sqrt{g^{L}}\right)}\right\|_{L^{2}(P)}\leq\frac{1}{2M_{l}^{3/2}}\left\|g^{L}-g^{U}\right\|_{L^{2}(P)}\leq\frac{\delta}{2M_{l}^{3/2}}$. Hence
$\leq\tilde{J}_{[]}\left(\delta,\mathcal{G}_{n},L^{2}(P)\right).$
Now we have shown that both the RHS of (23) and (24) are upper bounded by
Note that $\delta_{n}\asymp h_{Q}(g_{n},g_{0})$, and $h_{Q}(g_{n},g_{0})$ depends on the distance between the sieved neural network space where $g_{n}$ belongs to and the anisotropic Besov space where $g_{0}$ belongs to. To get $h_{Q}(g_{n},g_{0})$, we use Proposition 2 in Suzuki and Nitanda (2021), where the neural network parameters used in Lemma 4 are specified in the following way,
[…]
where $\epsilon=N^{-\tilde{\beta}}\log(N)^{-1}$, $m$ is a positive integer such that $0<\bar{\beta}<\min(m,m-1+1/s)$, $\delta=(1/s-1/r)_{+}$ (we use the $L^{2}(\mathcal{X})$ norm in (26) and let $r=2$), $\nu=(\tilde{\beta}-\delta)/(2\delta)$, and $U_{0}:=6dm(m+2)+2d$. $c_{(d.m)}$ is a constant that only depends on $d$ and $m$. The parameter $N$ controls the total number of neural network parameters, and we will allow $N$ to grow with the sample size $N=N(n)$ as $n\to\infty$. Hence we abbreviate the notations $\left\{L_{N},U_{N},S_{N},B_{N}\right\}$ to $\left\{L_{n},U_{n},S_{n},B_{n}\right\}$ in the following discussion. Proposition 2 in Suzuki and Nitanda (2021) states that, under (25),
Note that ()(*) is based on (26) and $g_{n}$ is taken to be the project of $g_{0}$ in $\mathcal{G}_{n}$.
  • (i)
    $r_{n}\sqrt{A_{n}+S_{n}+S_{n}\log(r_{n})}\lesssim\sqrt{n}$;
  • (ii)
    $\frac{1}{r_{n}}\gtrsim N^{-\tilde{\beta}}$.
We can simplify the first inequality (i) in the order of $N$ as follows
$=O(N\left(\log N\right)^{3}).$
Now the proof for the estimation error is complete,
[…]
Appendix D Additional Numeric Experiments Results
Figure 7: 900 generated images by DDIM with $r(X_{q})$ smaller than the leaset $r(X_{p})$ value. There are a total of 157779 out of 160000 such samples.
@@ Appendix D Additional Numeric Experiments Results
Clostridia 0.505
Coriobacteriia 0.433
Opitutae -0.396
Erysipelotrichi 0.392
Bacteroidia 0.382
Deinococci -0.371
Elusimicrobia -0.366
Synergistia -0.356
Acidimicrobiia -0.343
Rubrobacteria -0.333

### v2

@@ Appendix D Output activation function and sensitivity analysis
Figure 11: CelebA-64: DDIM evaluated on full training data using bounded softplus as the output activation function.
CelebA. The main results in Section 4.2 are based on the bounded sigmoid function with $\alpha=0.1$. Figure 11 presents the results when using the bounded softplus function, an asymmetric function, as the output activation. Additionally, we also included the logistic regression in Table 11(c) for the downstream feature detection. Although the numerical values of $r(X_{p}),r(X_{q})$ differ from the main analysis in Section 4.2, the general trend is the same: high concentration of homogeneous patterns of face images (left panel in Figure 11(e)), with $r(X_{q})$ concentrating around 0, shows possible model hallucinations; the most easily learned feature by DDIM is still the wearing_hat feature, and the most difficult one is the 5_o_Clock_Shadow. The major difference from Section 4.2 is that the evaluated $r(X_{p}),r(X_{q})$ cannot reach the upper bound 2, due to the asymmetric activation function.
Appendix E Additional Numerical Experiments
This section provides additional results for the numerical experiments in Section 4.
For the CelebA-64 data, Figure 12 shows 900 images that have small $r(X_{q})$ values around 0 and display a homogeneous pattern.
For the American Gut Project microbiome data, Figure 13 shows the alpha diversity plots in terms of the Shannon diversity and Gini-Simpson diversity. Figure 14 provides the sunburst plot on each level of the taxonomy table, where we conduct correlation analysis between the CLR-transformed relative abundance at each taxonomic level with the $r(X)$ evaluated in both observed and synthetic data. Positive correlation indicates the corresponding type of taxa is more abundant in the observed data, and negative correlation indicates more abundance in the synthetic data. Table 2 provides the numeric value of the correlation coefficients on the class level, with the top 10 classes of microbiome ranked by the absolute value of the correlation coefficient.
Figure 12: 900 generated images by DDIM with $r(X_{q})$ smaller than the leaset $r(X_{p})$ value. There are a total of 159719 out of 160000 such samples.
@@ Appendix E Additional Numerical Experiments
Clostridia 0.477
Coriobacteriia 0.421
Bacteroidia 0.404
Opitutae -0.389
Erysipelotrichi 0.376
Elusimicrobia -0.363
Spartobacteria -0.357
Synergistia -0.357
Deinococci -0.354
Acidimicrobiia -0.340
@@
Appendix F Simulation with Known Densities
Although the goal of this work is not to propose a new density ratio estimation method, it is still necessary to check the point estimation accuracy of the neural network estimator on known densities and compare its performance with existing density ratio estimation methods. In this simulation study, we start with a 2D case and extend to four high-dimensional cases with a latent 2D manifold.
In the 2D case, we simulate $P$ and $Q$ from mixtures of 3 normal densities, as shown in the first plot in Figure 15. The two distributions $P$ and $Q$ are defined as three-component Gaussian mixtures with shared mixture weights $\bm{w}=(w_{1},w_{2},w_{3})=(0.3,0.3,0.4)$:
$P(x)$ $=\sum_{k=1}^{3}w_{k}\,\mathcal{N}\!\left(x\mid\bm{\mu}_{k}^{(P)},\,\Sigma_{k}^{(P)}\right),$ (28)
$Q(x)$ $=\sum_{k=1}^{3}w_{k}\,\mathcal{N}\!\left(x\mid\bm{\mu}_{k}^{(Q)},\,\Sigma_{k}^{(Q)}\right).$ (29)
The 3-component mixture of normal densities represents 3 cases of distributional differences as shown in the first scatterplot in Figure 15: (1) the top-left group is the component where $P$ and $Q$ have the same density; (2) the top-right group is when $P$ and $Q$ have the same mean, but the covariance is different; (3) the bottom-left group is when $P$ and $Q$ have the same covariance structure, but the means are shifted. The numerical values of the normal parameters are set up as follows:
  1. (1)
    Top-left group (identical component). Both $P$ and $Q$ share the same mean and covariance:
    $\bm{\mu}_{1}^{(P)}=\bm{\mu}_{1}^{(Q)}=(-1,\,5)^{\!\top},\qquad\Sigma_{1}^{(P)}=\Sigma_{1}^{(Q)}=\begin{pmatrix}1&0\\[2.0pt] 0&1\end{pmatrix}$.
  2. (2)
    Top-right group (different covariance, same mean). Both distributions share the same mean but have different covariance structures:
    $\bm{\mu}_{2}^{(P)}=\bm{\mu}_{2}^{(Q)}=(5,\,5)^{\!\top},\qquad\Sigma_{2}^{(P)}=2\begin{pmatrix}1&-0.9\\[2.0pt] -0.9&1\end{pmatrix},\qquad\Sigma_{2}^{(Q)}=2\begin{pmatrix}1&0\\[2.0pt] 0&1\end{pmatrix}$.
  3. (3)
    Bottom-left group (mean shift, same covariance). The two distributions share the same covariance matrix but have shifted means:
    $\bm{\mu}_{3}^{(P)}=(-2,\,-2)^{\!\top},\qquad\bm{\mu}_{3}^{(Q)}=(0,\,0)^{\!\top},\qquad\Sigma_{3}^{(P)}=\Sigma_{3}^{(Q)}=\begin{pmatrix}1&0.5\\[2.0pt] 0.5&1\end{pmatrix}$.
For the high-dimensional simulation, we assume that there exists 2D low-dimensional structure to the high-dimensional densities. Let $d$ be the dimension of the high-dimensional densities, and $V_{p}$ follows the 3-component normal mixture distribution (28), and $V_{q}$ follows (29). To project the low-dimensional $V_{p},V_{q}$ onto $\mathbb{R}^{d}$, we first construct an orthonormal matrix $K\in\mathbb{R}^{d\times 2}$, whose columns form an orthonormal basis in $\mathbb{R}^{d}$. Specifically, we generate a random Gaussian matrix $G\in\mathbb{R}^{D\times 2}$ with entries $G_{ij}\sim\mathcal{N}(0,1)$ and obtain $K$ from the thin QR decomposition $G=QR,K=Q$, so that $K^{\top}K=I_{2}$. Then the low-dimensional sample $V_{p},V_{q}$ are projected to $\mathbb{R}^{d}$ using
$X_{i}=K\,V_{i}+\varepsilon_{i},\qquad\varepsilon_{i}\sim\mathcal{N}\!\left(0,\,\sigma^{2}I_{d}\right),$
and the resulting samples $X_{p},X_{q}$ are in $\mathbb{R}^{d}$, and their relative density ratio can be analytically computed as the theoretical truth. We choose to use 2D mixtures as the latent distributions simply because it is easy to visualize the performance.
We include three existing methods in this simulation comparison.
  1. 1.
    Unconstrained Least-Squares Importance Fitting (kanamori2009least, uLSIF) aims to directly minimize the loss function $\min_{r_{\theta}}\frac{1}{2}\mathbb{E}_{Q}\left\{\left(r_{\theta}(X)-r(X)\right)^{2}\right\}$ over a candidate of class of $r_{\theta}$ that are expressed as kernel expansions. This method assumes the true ratio function is within the reproducing kernel Hilbert space (RKHS) constructed by the kernel basis, and hence requires a global smooth structure of the true ratio function. We used the python package densratio (https://github.com/hoxo-m/densratio_py) and its implementation of uLSIF, and set the kernel bandwidth to be searched in $(0.1,0.3,1.0)$ by cross-validation.
  2. 2.
    Kullback–Leibler Importance Estimation Procedure (KLIEP) (sugiyama2007direct, KLIEP) also uses kernel basis to minimize the KL divergence $\min_{r_{\theta}}\text{D}_{\text{KL}}(p\|r_{\theta}q)$, and we use the adapt python package de2021adapt (https://adapt-python.github.io/adapt/). We use the RBF kernel, with at most 100 kernel centers, and use 5-fold cross-validation to automatically select the optimal kernel bandwidth.
  3. 3.
    The density ratio trick (DR-trick) via binary classification. This method is discussed in Section 2.3. We use logistic regression to estimate the binary classification probability $P(Z=1|X)$ in Section 2.3, and use the one-to-one mapping through the Bayes rule to recover the relative density ratio.
We denote our neural network model as MLP in Table 3. For each case shown in Table 3, we sample $n_{p}=n_{q}=1000$ equally from $P$ and $Q$ as the training sample, and another 1000 pairs as the testing sample. For each method, we use the training sample to train the relative density ratio function, and evaluate RDR on the testing sample and compute the testing MSE. For the neural network model based on MLP structure, to avoid overfit, we collect another 500 pairs of validation sample and stop the training process early if the validation loss does not improve by $10^{-5}$ over 5 consecutive epochs. The number of samples and all the tuning parameters in every method stay the same throughout all the 5 cases in Table 3. Note that including MLP, all of the methods in comparison can be optimized to achieve better performance for each case; however, our goal here is to use a simple simulation setting to study what might happen when extending these methods in high-dimensional RDR estimation.
Dim Noise MLP uLSIF KLIEP DR-trick
2 0.0 0.06 (0.04) 0.15 (0.03) 0.05 (0.01) 0.20 (0.01)
20 0.1 0.12 (0.03) 0.16 (0.04) 0.16 (0.01) 0.30 (0.01)
40 0.1 0.15 (0.02) 0.20 (0.04) 0.26 (0.02) 0.31 (0.01)
40 0.3 0.17 (0.02) 0.94 (0.03) 0.24 (0.02) 0.29 (0.01)
100 0.3 0.19 (0.01) 1.30 (0.02) 0.31 (0.01) 0.30 (0.01)
Table 3: Simulation comparison with existing methods. Each entry represents the mean (sd) of test MSE over 100 replicated studies.
Table 3 shows that MLP gives the second best performance in the 2D case, and the best performance for all other high-dimensional cases. To further understand the pros and cons of different methods, we provide visualizations in Figure 15 for all the 5 cases. The 2D plots are directly based on the testing sample. For the high-dimensional cases, we visualize the RDR values on the low-dimensional testing sample $V_{p},V_{q}$.
Based on Figure 15, the DR-trick cannot generalize well to very high-dimensions, since it tends to give very large or very small binary classifications probabilities to a few more extreme observations. uLSIF is a global kernel method that even in low-dimensions, it cannot correctly reflect the true RDR pattern, and it is most sensitive to the high-dimensional noise perturbation. KLIEP relies on the number of kernel centers; with the limit of such centers set at 100, as the dimension increases, the fixed number of kernel centers fails to capture the high-dimensional complexity of RDR, and the evaluated RDR tends to be more uniform around 1. For the 4-layer MLP, we also only have a fixed number of neural network parameters as the dimension increases, and inevitably, the performance deteriorates as the dimension increases. But the Hellinger-loss-based MLP is still able to capture the general pattern of the low-dimensional manifold especially in the mean-shift case. The caveat is that, when there is noise perturbation in high-dimensions, MLP tends to polarize the points that are supposed to have equal mass in $P$ and $Q$. This is because the noise perturbation in high-dimensions can shift the empirical support of the two samples and make certain points fall out of the overlapping support.
Figure 15: Simulation visualizations for each case shown in Table 3.
Figure 16: Simulation visualizations for each case (continued).